# Less is More: Compact Clue Selection for Efficient Retrieval-Augmented Generation Reasoning

### v6

@@ Abstract.
Current RAG retrievers are designed primarily for human readers, emphasizing complete, readable, and coherent paragraphs. However, LLMs benefit more from precise, compact, and well-structured input, which enhances reasoning quality and efficiency. Existing methods often rely on reranking or summarization to identify key sentences, but may suffer from semantic breaks and unfaithfulness. Thus, efficiently extracting and organizing answer-relevant clues from large-scale documents while reducing LLM reasoning costs remains a challenge for RAG. Inspired by Occam’s razor, we frame LLM-centric retrieval as a MinMax optimization: maximizing the extraction of potential clues and reranking them for well-organization, while minimizing reasoning costs by truncating to the smallest sufficient clues set. In this paper, we propose CompSelect, a Compact clue Selection mechanism for LLM-centric RAG, consisting of a clue extractor, a reranker, and a truncator. (1) The clue extractor first uses answer-containing sentences as fine-tuning targets, aiming to extract sufficient potential clues; (2) The reranker is trained to prioritize effective clues based on real LLM feedback; (3) The truncator uses the truncated text containing the minimum sufficient clues for answering the question as fine-tuning targets, thereby enabling efficient RAG reasoning. Experiments on three QA datasets show that CompSelect improves QA performance by approximately 11% and reduces Total Latency and Online Latency by approximately 17% and 67% compared to various baseline methods on both LLaMA3 and Qwen3. Further analysis confirms its robustness to unreliable retrieval and generalization across different scenarios, offering a scalable and cost-efficient solution for web-scale RAG applications.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06
@@ 1. Introduction
Figure 1. SubEM performance and latency on the NQ test set for different sizes of the LLaMA3 series, including the reranking method RECOMP-extr and the abstractive method Refiner.
@@
Retrieval-augmented generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Zhou et al., 2024) has emerged as a powerful paradigm to enhance large language models (LLMs) with external knowledge on various tasks (Chen et al., 2022; Huang et al., 2023; Gao et al., 2023). However, the retrievers currently used in RAG are designed primarily for human readers, favoring long and coherent passages to enhance readability. This paradigm is less effective when the downstream consumer is an LLM (Sauchuk et al., 2022), which typically reasons better when given precise, compact, and well-structured inputs. As shown in Figure 1, we compare the Top-1 document, the full set of retrieved documents (approximately 5), and the answer-containing sentences (Oracle Clues) on the NQ dataset across different sizes of LLaMA3. The results indicate that, regardless of whether it is a 1B model or the larger 70B model, structuring retrieved documents into a concise and efficient manner (as oracle clues) will lead to better reasoning performance and lower system latency. These findings highlight the need for compact clue selection mechanisms tailored to LLM reasoning in RAG systems.
@@
Prior work has explored techniques such as reranking (Xu et al., 2024b; Ke et al., 2024; Qin et al., 2024) and summarization (Hwang et al., 2025; Zhu et al., 2024a) to help RAG identify key sentences. The former reranks the sentences from retrieved documents based on metrics such as answer contribution or user preference (Mao et al., 2016; Zhu et al., 2023). The latter identifies query-relevant sentences either through extractive selection or abstractive generation. Each approach has limitations: sentence-level reranking may disrupt the semantic structure of documents; extractive methods are limited in integrating dispersed information and capturing global context; and abstractive methods increase inference costs and may alter the original documents, thereby introducing issues of unfaithfulness. As shown in Figure 1, the reranking method RECOMP-extr (Xu et al., 2024b) primarily reduces reasoning latency but sacrifices QA performance, whereas the abstractive method Refiner (Li et al., 2024) improves QA performance but increases system latency compared to baseline models. Therefore, efficiently extracting and organizing answer-relevant clues from large-scale documents while reducing LLM reasoning overhead remains an open challenge for RAG.
@@
Inspired by Occam’s razor (d’Ockham, 1938; Rasmussen and Ghahramani, 2000), we frame LLM-centric retrieval as a MinMax optimization problem. We first leverage contextual information to identify potential clues, forming the maximal subset capable of answering the query. Next, we rerank these clues based on their completeness and relevance to the query, moving the most effective ones at the forefront. Finally, we truncate the set to retain only the essential clues, thereby minimizing the context required for RAG. As shown in Figure 2, our method highlights potential clues (red background), reranks them, and surfaces the correct answer clue at the top, while filtering out redundant ones to improve the information density of reasoning clues for RAG.
@@
Experiments on NQ, TriviaQA, and HotpotQA datasets show that CompSelect consistently outperforms baselines while requiring significantly less context for inference on both LLaMA3 and Qwen3. In addition to performance gains, it proves robust to unreliable retrieval and generalizes well across scenarios. By aligning retrieval with the reasoning needs of LLMs, CompSelect provides a scalable, cost-efficient solution for web-scale RAG applications  [^1] Our code is available at https://anonymous.4open.science/r/CompSelect-E463/.
@@
  • We frame an LLM-centered compact clue selection mechanism as a MinMax optimization, motivated by our finding that precise, compact, and well-organized inputs enhance both the LLM’s reasoning performance and efficiency.
@@
  • Experiments on three datasets demonstrate that compact clue selection not only enhances inference performance and reduces system latency, but also provides robustness against unreliable retrieval and supports generalization across tasks.
Figure 3. The architecture of CompSelect consists of three modules: a clue extractor, a reranker, and a truncator. The top shows their training strategies and annotated data, and the bottom illustrates the inference process of compact clue selection.
@@ 2. Related Work
Reranking methods prioritize the most relevant content from retrieved documents. Classical methods like BM25 (Robertson et al., 2009) rank documents based on term frequency, inverse document frequency, and document length. Neural rerankers (Mitra et al., 2018; Mitra and Craswell, 2017), such as BGE-Reranker (Xiao et al., 2024) and RECOMP-extr (Xu et al., 2024b), use dense embeddings or fine-tuned cross-encoders to select salient sentences. RichRAG (Wang et al., 2025) employs a generative list-wise ranker to produce and rank candidate documents, while  Ke et al. (2024) introduces a bridge mechanism to better connect retrievers and LLMs. CPC (Liskavets et al., 2025) ranks sentences with context-aware embeddings. Although reranking methods maintain low latency, they may harm RAG performance by disrupting the original logical structure of the document and can generate unfaithful clues.
Extractive methods aim to select key information from retrieved documents and can be further categorized into fragment selection and token pruning. Fragment selection methods, such as EXIT (Hwang et al., 2025), use sentence-level classifiers to pick the most relevant sentences, achieving low latency but often relying on rigid selection criteria that may not adapt well to query complexity or document quality. Token-pruning methods, including LongLLMLingua (Jiang et al., 2024) and Selective-Context (Li et al., 2023), leverage LLMs to compute the informativeness of each token and remove low-information tokens, which may cause context fragmentation and higher latency. Other approaches, such as AdaComp (Zhang et al., 2024), BlendFilter (Wang et al., 2024), and EviOmni (Zhao et al., 2025), employ adaptive selection, knowledge blending, or reinforcement learning to improve reasoning efficiency and accuracy.
Abstractive methods generate query-relevant content via autoregressive generation. RECOMP-abs (Xu et al., 2024b) uses a T5-based model (Raffel et al., 2020), Refiner (Li et al., 2024) hierarchically structures query-relevant content with larger LLMs, and BottleNeck (Zhu et al., 2024a) uses reinforcement learning with information bottleneck theory to generate task-relevant summaries. FILCO (Wang et al., 2023) trains a filtering model to dynamically select key sentences and jointly learns with the generator for end-to-end distillation. SEER (Zhao et al., 2024) uses self-alignment to select evidence via expert evaluation, but still depends on domain-specific experts and can be sensitive to noisy retrieval. These approaches reorganize scattered information into concise outputs, and as a result of their autoregressive generation, they typically incur higher latency.
In summary, reranking methods improve efficiency but reduce QA accuracy, extractive methods balance efficiency and fidelity with risk of fragmentation, and abstractive methods flexibly restructure content at higher costs. In contrast, our framework maximizes sufficiency, prioritizes effectiveness, and minimizes reasoning costs, enhancing both efficiency and accuracy in LLM-centric RAG.
@@ 3. Problem Formulation
Given a query $q$ and a set of retrieved documents $\mathcal{D}=\{d_{1},\ldots,d_{n}\}$, where each document $d_{i}$ consists of a set of sentences $\mathcal{S}_{i}=\{s^{i}_{1},\ldots,s^{i}_{n_{i}}\}$, $n_{i}$} is the number of sentences in $d_{i}$. The objective of the noise filtering task is to identify an optimal subset $\mathcal{S}^{*}\subseteq\bigcup_{i=1}^{n}\mathcal{S}_{i}$ such that a language model $f_{\theta}$ generates the correct answer $y$ for the query $q$ with the highest probability. The optimal subset $\mathcal{S}^{*}$ can be determined by the following MinMax optimization:
(1) $\mathcal{S}^{*}$ $=\arg\min\limits_{S^{\prime}\in\mathcal{S}_{\text{max}}}|S^{\prime}|,$
(2) $\mathcal{S}_{\text{max}}$ $=\arg\max\limits_{S\subseteq\bigcup_{i=1}^{n}S_{i}}\log P_{\theta}(y\mid S,q),$
where $\mathcal{S^{{}^{\prime}}}$ is the subset that is most capable of producing the correct answer, and $|\mathcal{S^{{}^{\prime}}}|$ is the number of sentences in $\mathcal{S^{{}^{\prime}}}$. The selection of $\mathcal{S}^{*}$ should dynamically adapt to the real feedback of a RAG system to balance informativeness and conciseness. The problem can be formalized as an NP-hard combinatorial optimization problem (Wu et al., 2023), selecting the smallest, most relevant answer clues from a large set of documents to improve answer accuracy.
@@ 4. Methodology
In this section, we introduce CompSelect, a three-stage framework for selecting and prioritizing relevant information in RAG, as illustrated in Figure 3. CompSelect consists of three modules: a clue extractor, a clue reranker, and an adaptive truncator. The clue extractor identifies potential answer clues from multiple documents, reducing the search space while retaining relevant information. The reranker then prioritizes the most relevant clues using a pairwise loss strategy. Finally, the adaptive truncator adaptively selects the minimal necessary context for each query, improving inference efficiency and answer accuracy.
@@ 4.1. Clue Extractor
The goal of clue extractor is to identify potential answer clues from multiple documents and construct a smaller query-relevant candidate set to reduce the search space. We compare the performance of answer-containing sentences and the original retrieved documents in downstream tasks, as shown in Figure 1. The results indicate that filtering out low-value information from documents benefits RAG reasoning. Although not all answer-containing sentences are query-relevant, they approximate the maximal subset capable of addressing user queries, and can serve as the optimization target for the clue extractor.
To guide the extraction of informative sentences, we first extract sentences from the retrieved document collection that contain the ground-truth answer as targets for extraction. Given a query $q$, the ground-truth answer $y$ and retrieved sentences $\mathcal{S}=\{s_{1},s_{2},\dots,s_{n}\}$, the answer-containing sentences are defined as:
@@
Finally, the clue extractor is capable of generating the potential candidate clues based on the user query and retrieved documents during inference:
@@
Table 1. Performance comparison of different strategies on NQ, TriviaQA, and HotpotQA. We use Llama3-8B-Instruct as the generator and SubEM for evaluation.
@@ 4.1.1. KNN-based Extraction
As shown in Table 1, selecting answer-containing sentences substantially improves performance on single-hop QA datasets such as NQ, raising the SubEM score from 45.87 to 56.73. This highlights the benefit of focusing on relevant context for simple queries. For multi-hop or more complex QA datasets, including TriviaQA and HotpotQA, incorporating answer-containing sentences also leads to performance gains. Moreover, further retrieving semantically similar sentences using a KNN-based strategy yields additional improvements, indicating that extra contextual information beyond strictly answer-containing sentences is beneficial for multi-hop reasoning and complex query understanding.
Based on these observations, we propose a KNN-based extraction strategy. For simple questions, we select answer-containing sentences as optimization targets to provide key information. For complex questions, we additionally retrieve semantically similar sentences to supplement context. By using both answer-containing and KNN-based similar sentences as optimization targets, the extractor adapts to queries of varying complexity, improving reasoning accuracy while maintaining a concise candidate set.
@@ 4.2.1. Training
We use the real RAG-generated feedback to annotate the training data for the reranker, as the QA performance on complex questions heavily depends on the characteristics of the generation module. First, we pair each of the extracted clue sentences $s_{j}^{c}\in\mathcal{S}^{c}$ with the query $q$ as $(s_{j}^{c},q)$, where sentence $s_{j}^{c}$ that enables the downstream generation module to produce the correct answer for $q$ is considered as positive sample $s_{\text{pos}}$, while other sentences are treated as negative samples $s_{\text{neg}}$ [^2] If no candidate clues can generate the correct answer, or if all samples can generate the correct answer, the sample will be removed from the annotated data.. The $Reranker$ aims to minimize the following pairwise loss function (Karpukhin et al., 2020) to improve relevance ranking:

### v7

@@ Abstract.
Current RAG retrievers are designed primarily for human readers, emphasizing complete, readable, and coherent paragraphs. However, Large Language Models (LLMs) benefit more from precise, compact, and well-structured input, which enhances reasoning quality and efficiency. Existing methods rely on reranking or summarization to identify key sentences, but may introduce semantic breaks and unfaithfulness. Thus, efficiently extracting and organizing answer-relevant clues from large-scale documents while reducing LLM reasoning costs remains challenging in RAG systems. Inspired by Occam’s razor, we frame LLM-centric retrieval as MinMax optimization: maximizing the extraction of potential clues and reranking them for well-organization, while minimizing reasoning costs by truncating to the smallest sufficient set of clues. In this paper, we propose CompSelect, a compact clue selection mechanism for LLM-centric RAG, consisting of a clue extractor, a reranker, and a truncator. (1) The clue extractor first uses answer-containing sentences as fine-tuning targets, aiming to extract sufficient potential clues; (2) The reranker is trained to prioritize effective clues based on real LLM feedback; (3) The truncator uses the truncated text containing the minimum sufficient clues for answering the question as fine-tuning targets, thereby enabling efficient RAG reasoning. Experiments on three QA datasets demonstrate that CompSelect improves performance while reducing both total and online latency compared to a range of baseline methods. Further analysis also confirms its robustness to unreliable retrieval and generalization across different scenarios.
Retrieval-Augmented Generation, Large Language Models, Information Retrieval
journalyear: 2026copyright: ccconference: Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, United Arab Emirates.booktitle: Proceedings of the ACM Web Conference 2026 (WWW ’26), April 13–17, 2026, Dubai, United Arab Emiratesdoi: 10.1145/3774904.3792158isbn: 979-8-4007-2307-0/2026/04ccs: Information systems Retrieval models and ranking
@@ 1. Introduction
Retrieval-Augmented Generation (RAG) (Lewis et al., 2020; Gao et al., 2023; Zhou et al., 2024) has emerged as a promising paradigm to enhance Large Language Models (LLMs) with external knowledge for various tasks (Chen et al., 2022; Huang et al., 2023; Chen et al., 2025b; Wang et al., 2025c). However, the retrievers currently used in RAG are designed primarily for human readers, favoring long and coherent passages to enhance readability. This paradigm is less effective when the downstream consumer is an LLM (Sauchuk et al., 2022), which typically performs better when provided with precise, compact, and well-structured inputs. As shown in Figure 1, we compare the Top-1 document, the full set of retrieved documents (approximately 5), and the answer-containing sentences (Oracle Clues) on the NQ dataset across different LLaMA3 model sizes. The results indicate that, regardless of whether it is a 1B model or the larger 70B model, structuring retrieved documents into a concise and efficient form (as Oracle Clues) consistently improves reasoning performance and reduces system latency. These findings highlight the need for compact clue selection mechanisms tailored to LLM reasoning in RAG systems.
@@
Prior work has explored techniques such as reranking (Xu et al., 2024a; Ke et al., 2024; Qin et al., 2024) and summarization (Hwang et al., 2025; Zhu et al., 2024b) to help RAG identify key sentences. The former reranks the sentences from retrieved documents based on metrics such as answer contribution or user preference (Mao et al., 2016; Zhu et al., 2023). The latter identifies query-relevant sentences either through extractive selection or abstractive generation. Each approach has limitations: sentence-level reranking may disrupt the semantic structure of documents; extractive methods are limited in integrating dispersed information and capturing global context; and abstractive methods increase inference costs and may alter the original documents, thereby introducing issues of unfaithfulness. As shown in Figure 1, the reranking method RECOMP-extr (Xu et al., 2024a) primarily reduces reasoning latency but sacrifices QA performance, whereas the abstractive method Refiner (Li et al., 2024) improves QA performance but increases system latency compared to baseline models. Therefore, efficiently extracting and organizing answer-relevant clues from large-scale documents while reducing LLM reasoning overhead remains an open challenge for RAG (Li et al., 2025; Su et al., 2025).
Figure 1. SubEM performance and latency on the NQ test set across different LLaMA3 model sizes, including the reranking method RECOMP-extr and the abstractive method Refiner.
@@
Inspired by Occam’s razor (d’Ockham, 1938; Rasmussen and Ghahramani, 2000), we frame LLM-centric retrieval as a MinMax optimization problem. We first leverage contextual information to identify potential clues, forming the maximal subset capable of answering the query. Next, we rerank these clues based on their completeness and relevance to the query, moving the most effective ones at the forefront. Finally, we truncate the set to retain only the essential clues, thereby minimizing the context required for RAG.
@@
Experiments on NQ, TriviaQA, and HotpotQA datasets show that CompSelect consistently outperforms baselines while requiring significantly less context for inference on both LLaMA3 and Qwen3. In addition to performance gains, it proves robust to unreliable retrieval and generalizes well across scenarios. By aligning retrieval with the reasoning needs of LLMs, CompSelect provides a scalable, cost-efficient solution for web-scale RAG applications.
@@
  • We frame the LLM-centric compact clue selection mechanism as MinMax optimization, motivated by our finding that precise, compact, and well-organized inputs enhance both the LLM’s reasoning performance and efficiency.
@@
  • Experiments on three datasets demonstrate that compact clue selection not only enhances inference performance and reduces system latency, but provides robustness against unreliable retrieval and supports generalization across tasks.
Figure 2. Architecture of CompSelect, consisting of three modules: a clue extractor, a reranker, and a truncator.
@@ 2. Related Work
Reranking methods prioritize the most relevant content from retrieved documents. Classical methods like BM25 (Robertson et al., 2009) rank documents based on term frequency, inverse document frequency, and document length. Neural rerankers (Mitra et al., 2018; Mitra and Craswell, 2017), such as BGE-Reranker (Xiao et al., 2024) and RECOMP-extr (Xu et al., 2024a), use dense embeddings or fine-tuned cross-encoders to select salient sentences. RichRAG (Wang et al., 2025a) employs a generative list-wise ranker to produce and rank candidate documents, while  Ke et al. (2024) introduces a bridge mechanism to better connect retrievers and LLMs. CPC (Liskavets et al., 2025) ranks sentences with context-aware embeddings. Although reranking methods maintain low latency, they may harm RAG performance by disrupting the original logical structure of the document and can generate unfaithful clues.
Extractive methods aim to select key information from retrieved documents and can be further categorized into fragment selection and token pruning. Fragment selection methods, such as EXIT (Hwang et al., 2025), use sentence-level classifiers to pick the most relevant sentences, achieving low latency but often relying on rigid selection criteria that may not adapt well to query complexity or document quality. Token-pruning methods, including LongLLMLingua (Jiang et al., 2024) and Selective-Context (Li et al., 2023), leverage LLMs to compute the informativeness of each token and remove low-information tokens, which may cause context fragmentation and higher latency. Other approaches, such as AdaComp (Zhang et al., 2024), BlendFilter (Wang et al., 2024), ClueAnchor (Chen et al., 2025a) and EviOmni (Zhao et al., 2025), employ adaptive selection, knowledge blending, or reinforcement learning to improve reasoning efficiency and accuracy.
Abstractive methods generate query-relevant content via autoregressive generation. RECOMP-abs (Xu et al., 2024a) uses a T5-based model (Raffel et al., 2020), Refiner (Li et al., 2024) hierarchically structures query-relevant content with larger LLMs, and BottleNeck (Zhu et al., 2024b) uses reinforcement learning with information bottleneck theory to generate task-relevant summaries. FILCO (Wang et al., 2023) trains a filtering model to dynamically select key sentences and jointly learns with the generator for end-to-end distillation. SEER (Zhao et al., 2024) uses self-alignment to select evidence via expert evaluation, but still depends on domain-specific experts and can be sensitive to noisy retrieval. These approaches reorganize scattered information into concise outputs, and as a result of their autoregressive generation, they typically incur higher latency.
In contrast, our work maximizes sufficiency, prioritizes effectiveness, and minimizes reasoning costs, enhancing both efficiency and accuracy in LLM-centric RAG.
@@ 3. Problem Formulation
Given a query $q$ and its retrieved document set $\mathcal{D}=\{d_{1},\ldots,d_{n}\}$, where each document $d_{i}$ consists of a set of sentences $\mathcal{S}_{i}=\{s^{i}_{1},\ldots,s^{i}_{n_{i}}\}$, with $n_{i}$ denoting the number of sentences in $d_{i}$. The objective of our task is to identify an optimal subset $\mathcal{S}^{*}\subseteq\bigcup_{i=1}^{n}\mathcal{S}_{i}$ such that the language model $f_{\theta}$ generates the correct answer $y$ for the query $q$, while minimizing the subset size. The optimal subset $\mathcal{S}^{*}$ can be determined by the following MinMax optimization:
(1) $\mathcal{S}^{*}$ $=\arg\min\limits_{\mathcal{S^{\prime}}\subseteq\mathcal{S}_{\text{max}}}|\mathcal{S^{\prime}}|,$
(2) $\mathcal{S}_{\text{max}}$ $=\arg\max\limits_{\mathcal{S}\subseteq\bigcup_{i=1}^{n}\mathcal{S}_{i}}\log P_{\theta}(y\mid\mathcal{S},q),$
where $\mathcal{S^{\prime}}$ is the minimal subset of $\mathcal{S}_{\max}$ that still allows the model $f_{\theta}$ to generate the correct answer and $|\mathcal{S^{\prime}}|$ denotes the number of sentences in $\mathcal{S^{\prime}}$. The selection of $\mathcal{S}^{*}$ should dynamically adapt to the real feedback of a RAG system to balance informativeness and conciseness.
@@ 4. Methodology
In this section, we introduce CompSelect, a three-stage framework for selecting and prioritizing relevant information in RAG, as illustrated in Figure 2. CompSelect consists of three modules: a clue extractor, a clue reranker, and an adaptive truncator. The clue extractor identifies candidate answer clues from multiple documents, reducing the search space while retaining relevant information. The reranker then prioritizes the most relevant clues using a pairwise loss strategy. Finally, the truncator adaptively selects the minimal necessary context for each query, thereby improving inference efficiency and answer accuracy  [^1] Our code is available at https://github.com/zqc1023/CompSelect..
@@ 4.1. Clue Extractor
The goal of the clue extractor is to identify potential answer clues from multiple documents and construct a smaller query-relevant candidate set to reduce the search space. We compare the performance of answer-containing sentences and the original retrieved documents in downstream tasks, as shown in Figure 1. The results indicate that filtering out low-value information from documents benefits RAG reasoning. Although not all answer-containing sentences are query-relevant, they approximate the maximal subset capable of addressing user queries, and can serve as the optimization target for the clue extractor.
To guide the extraction of informative sentences, we first identify sentences from the retrieved documents that contain the ground-truth answer as targets for extraction. Given a query $q$ and the ground-truth answer $y$, we denote by $\mathcal{S}=\{s_{1},\dots,s_{n}\}$ the set of sentences appearing in the retrieved documents, with sentences preserving their original order within and across documents. The answer-containing sentences are defined as:
@@
Finally, the clue extractor is capable of generating the potential candidate clues based on the user query and retrieved sentences during inference:
@@
Table 1. SubEM (%) comparison of different strategies on NQ, TriviaQA, and HotpotQA test set. We use LLaMA3-8B-Instruct as the generator. $\epsilon$ shows the KNN threshold and higher values introduce more contextual sentences.
@@ 4.1.1. KNN-based Extraction
We further augment the clue extractor with a KNN-based strategy. As shown in Table 1, selecting answer-containing sentences substantially improves performance on single-hop QA datasets such as NQ, raising the SubEM score from 45.87 to 56.73. This highlights the benefit of focusing on relevant context for simple queries. For multi-hop or more complex QA datasets (e.g., TriviaQA and HotpotQA), incorporating answer-containing sentences also leads to performance gains. Moreover, further retrieving semantically similar sentences using a KNN-based strategy yields additional improvements, indicating that extra contextual information beyond strictly answer-containing sentences is beneficial for multi-hop reasoning and complex query understanding. Based on these observations, we propose a KNN-based extraction strategy. For simple questions, we select answer-containing sentences as optimization targets to provide key information. For complex questions, we additionally retrieve semantically similar sentences to supplement context. By using both answer-containing and KNN-based similar sentences as optimization targets, the extractor adapts to queries of varying complexity, improving reasoning accuracy while maintaining a concise candidate set.
@@ 4.2.1. Training
We use real downstream generator feedback from the RAG system to annotate the training data for the reranker, as QA performance on complex questions heavily depends on the characteristics of the generation module. First, we pair each of the extracted clue sentences $s_{j}^{c}\in\mathcal{S}^{c}$ with the query $q$ as $(s_{j}^{c},q)$, where sentence $s_{j}^{c}$ that enables the downstream generation module to produce the correct answer for $q$ is considered as positive sample $s_{\text{pos}}$, while other sentences are treated as negative samples $s_{\text{neg}}$ [^2] If no candidate clues can generate the correct answer, or if all samples can generate the correct answer, the sample will be removed from the annotated data.. The $Reranker$ aims to minimize the following pairwise loss function (Karpukhin et al., 2020) to improve relevance ranking: